OpenStudy (anonymous):

i need help with 24 + 8(x – 7) = 4(x + 6) + 4x

hartnn (hartnn):

use this: a(b+c)=ab+ac so whats 8(x-7) = ? and 4(x+6)= ?

OpenStudy (anonymous):

24 + 8x - 56 = 4x + 24 + 4x , Can you do next step ? try it ...then i'll help you

OpenStudy (anonymous):

i got 8x-56 and 4x+24

OpenStudy (anonymous):

i did normal distribuitve property, but none of the answers of the multiple choice came up

hartnn (hartnn):

very good,so now u have 24 + 8x - 56 = 4x + 24 + 4x what u got as final answer?

OpenStudy (anonymous):

i got 56

hartnn (hartnn):

is there a option called 'no value of x' or no solution ?

OpenStudy (anonymous):

think so, it no solution

OpenStudy (anonymous):

yes, why is there no solution?!?!

hartnn (hartnn):

okk,since 8x from both sides cancel out...

OpenStudy (anonymous):

becuz not ever function has 56 = 0 , never

OpenStudy (anonymous):

how 56 = 0 ? cannot , that's why we have no solution

hartnn (hartnn):

so now u don't have any x to calculate from.

OpenStudy (anonymous):

ohhhhhh thanks, cant believe i didnt see that

hartnn (hartnn):

welcome :)